# Parameter estimation with uncertainty quantification from continuous measurement data using neural network ensembles

### v1

@@ Abstract
We show that ensembles of deep neural networks, called deep ensembles, can be used to perform quantum parameter estimation while also providing a means for quantifying uncertainty in parameter estimates, which is a key advantage of using Bayesian inference for parameter estimation. These models are shown to be more robust to noise in the measurement results used to perform the parameter estimation as well as noise in the data used to train them. We also show that much less data is needed to achieve comparable performance to Bayesian inference based estimation, which is known to reach the ultimate precision limit as more data is collected, than was used in previous proposals.
@@ I Introduction
Firstly Bayesian inference protocols generally require derivation of an analytic expression of the likelihood distribution that the measurement data is assumed to be generated from. This is not always feasible for many systems under ideal conditions and for simple systems this can become intractable when one attempts to model the affects of experimentally imperfections. This may be circumvented by use of so-called ‘likelihood-free’ methods such as approximate Bayesian computation (ABC) [12, 13] but this method presents it’s own challenges such as a drastic increase in sample complexity when the number of parameters being estimated increases, requiring access to a simulator which can generate samples from the likelihood distribution, and selection of a suitable discrepancy measure (and summary statistic) for the rejection sampling routine [14, 15, 12].
@@
Figure 1: Diagram of quantum parameter estimation procedure using continuous photon number measurement of a two-level system (qubit) coupled to an external bath (the environment) with coupling strength $\gamma$ and continuously driven by an external laser field.
Recently approaches based on the use of machine learning (ML) algorithms, in particular neural networks (NNs), have gained popularity for performing quantum parameter estimation [18, 19, 20, 21, 22, 23]. In contrast to Bayesian inference, NN based approaches provide faster inference time, owing to their highly parallelizable inference process, while retaining the benefits of likelihood-free methods like ABC as they can learn complex functions from input-output pairs of training examples due to their high expressive power [24]. Furthermore, the ABC-based estimation method presented in [13] requires the a priori preparation of multiple sets of measurement results for different values of $\theta$ to perform inference. While an ML approach also requires this for the initial model training phase it is typically no longer needed at inference time unlike the ABC method. In Ref. [21] it was shown that a NN can be trained to perform parameter estimation given input-output pairs of quantum measurement results from a continuous measurement of a quantum probe system and the corresponding parameter values which generated those measurement results. This approach however utilized a large amount of training data and only provides point estimates without uncertainty quantification. Here we extend this approach to provide uncertainty estimates, thereby retaining a benefit of Bayesian approaches, and also drastically reducing the amount of training data required to train the model. We also study the robustness of our method to different types of noise in the training data and measurement results provided to the model at inference time.
@@ III.1 Deep ensembles
Model ensembling is a simple but powerful method in machine learning. Ensemble methods employ multiple models, typically ‘weak learners’ as individually these models have suboptimal performance, to create a single overall model by aggregating the predictions of the component models. The resulting model often has better performance, in terms of both bias and variance, than the individual models in isolation [28]. A state-of-the-art method for predictive uncertainty quantification for deep learning based models are deep ensembles [29, 30, 17]. For regression problems, such as the one considered in this work, deep ensembles work by training $M$ independent NNs to learn a continuous probability distribution over predictions. The overall model is then treated as a mixture of these distributions.
Deep ensembles are preferable to Bayesian NNs (BNNs) [14], which are another common tool for quantifying uncertainty in deep learning, for two reasons. Firstly they require minimal changes to the standard NN training process and their training is easily scalable via parallelization since each network in the ensemble can be trained completely independently. Secondly BNNs face the same computational problem as Bayesian inference in general i.e. the intractability of computing the posterior distribution over model parameters exactly [14, 29]. In fact due to the large number of parameters in NNs standard Bayesian sampling methods like MCMC are often too computationally expensive to use for BNN training and methods such as variation inference, which are faster but have weaker theoretical guarantees compared to MCMC, must be used [17].
@@ III.2 Model training
respectively [29]. Note that the distributions need not be Gaussian, the negative log likelihood of any probability distribution may be used if deemed appropriate and one is able to compute the relevant moments.
@@ III.3 Hyperparameter Tuning
Ref. [21] trained single NNs on a dataset of $4\times 10^{6}$ trajectory-parameter pairs [32]. With the use of hyperparameter tuning we are able to achieve the same level of performance with only $1\%$ ($4\times 10^{4}$) of the original training data. This is also much lower than the amount of data needed for ABC which used 1000 trajectories generated for 100 different values of $\Delta$ which results in a total dataset size of $10^{5}$ [13, 33]. Note also that the ABC approach used longer trajectories with $N=200$ time delay measurements as opposed to the $N=48$ used in this work. We split the training data into training and validation sets with a $80\%$ and $20\%$ split respectively. We use the tree-structured Parzen estimator optimization algorithm provided by the Optuna framework to perform hyperparameter tuning of the single network models presented in this work [34]. We utilize both trail pruning and early stopping during our tuning and training processes respectively and tune for 20 trials. For the single models the loss function used during training was also tuned as part of the hyperparameter tuning process with the possible choices being the root mean squared error (RMSE) loss or the mean squared logarithmic error (MSLE) loss [21]. Further details on the tuned hyperparemeters can be found in Appendix B. For experiment tracking, validation metric logging, and model storage we utilize the open source framework MLflow [35]. We make the MLflow logs which contain learning curve plots, model weights, and hyperparameter settings for all the models trained in the present work available at [36].
III.4 Training noise robustness
@@ III.5.2 Uncertainty quantification under data shift
(b) Average predicted variance $\sigma^{2}$ by deep ensemble for test data generated from the training distribution which includes no time jitter and test data with noise jitter following the Gaussian distribution $\mathcal{N}(0,\sigma_{\tau})$.
@@
An important property to examine for models which provide uncertainty estimates is how the predicted uncertainty changes when the model is provided out-of-distribution examples e.g. in the case of data shift [29]. Ideally we would want our model to have high predictive uncertainty when given input data from a distribution that is far from the training distribution i.e. we would expect the predicted Gaussian mixture to have greater variance $\sigma^{2}$. Fig. 7(a) shows the average predicted uncertainty of the deep ensemble model when given sets of time delays from systems in which the Rabi frequency differs from the value used to generate the training data. We can see that as the Rabi frequency deviates further from $\Omega=1$ the predictive uncertainty of our model increases. The same trend can be observed in Fig. 7(b) which shows the average predicted uncertainty for varying levels of time jitter noise. This can be useful as a way of detecting imperfections or miscalibration in the experiential setup used to generate the input time delay results.
IV Computational Efficiency
@@
(b) Average inference times over 10 runs for Bayesian inference and quantized deep ensemble for various numbers of trajectories.
@@
We also compare inference times by measuring the wall clock time of both the quantized deep ensemble and an optimized Hamiltonian Monte Carlo based Bayesian inference routine for estimating $\Delta$ given delay measurements from different numbers of trajectories. We use the state-of-the-art No U-Turn Sampler (NUTS) provided by the PyMC framework as our specific HMC method [44, 45]. For inference with NUTS we use 4 chains and draw 1,000 posterior samples per chain which we parallelize over 4 cores. The average inference time for the two methods over 20 different runs of the inference on an AMD Ryzen 9 3900X 12-Core Processor are shown in Fig. 8(b). We can see that the highly parallelizable nature of neural network inference, i.e. batched inference, allows for orders of magnitude faster inference speed over the NUTS method despite the need to perform a forward pass for each of the 10 NNs in the deep ensemble.
V Conclusion
We have proposed a method for using NNs to perform parameter estimation which provides uncertainty estimates in addition to point estimates of the parameters. Since our method is based on NNs it is easily generalizable to higher dimensional estimation problems by simply adding additional neurons in the output layer of the network. Furthermore, our scheme is easily generalizable to parameter estimation settings in which the measurement results are not independent and identically distributed, unlike the case of the TLS studied in the present work, where recurrent NNs become more appropriate to use over the permutation-invariant architectures discussed in this work.
We leave for future work the exploration of other NN architectures, such as deep sets, that possess permutation invariance and the ability to naturally handle inputs of varying length [46]. In this work we performed hyperparameter tuning on a single model and used the optimal hyperparameters from that search to train the deep ensemble due to time constraints. An interesting question to answer would be if hyperparameter tuning on the ensemble level, while being more computationally expensive, could deliver even better results. Lastly in this work we considered ensembles in which all models have the same architecture but this is by no means a necessity. Further exploration of how using ensembles of networks with different architectures, e.g. different activation functions and numbers of hidden layers, would impact ensemble performance on the parameter estimation task would be useful.

### v2

# arXiv:2509.10756v2[quant-ph] 01 Mar 2026

@@ Abstract
We show that ensembles of deep neural networks, called deep ensembles, can be used to perform quantum parameter estimation while also providing a means for quantifying uncertainty in parameter estimates, which is a key advantage of using Bayesian inference for parameter estimation that is lost when using existing machine learning methods. We show that optimizing for both accurate parameter estimates and well calibrated uncertainty estimates does not lead to degradation in the former as opposed to only optimizing for accuracy. We also show that the drift detection capabilities of these ensemble models can be used to detect drift in the experimental data used during inference. These results suggest that such models could enable accurate, real-time parameter estimation with quantified uncertainty, making them promising candidates for deployment in experimental settings.
@@ I Introduction
Firstly Bayesian inference protocols generally require derivation of an analytic expression of the likelihood distribution that the measurement data is assumed to be generated from. This is not always feasible for many systems under ideal conditions and for simple systems this can become intractable when one attempts to model the effects of experimentally imperfections. This may be circumvented by use of so-called ‘likelihood-free’ methods such as approximate Bayesian computation (ABC) [46, 10] but this method presents it’s own challenges such as a drastic increase in sample complexity when the number of parameters being estimated increases, requiring access to a simulator which can generate samples from the likelihood distribution, and selection of a suitable discrepancy measure (and summary statistic) for the rejection sampling routine [36, 18, 46].
@@
Figure 1: Diagram of quantum parameter estimation procedure using continuous photon counting measurement of a two-level system (qubit) coupled to an external bath (the environment) with coupling strength $\gamma$ and continuously driven by an external laser field.
Recently approaches based on the use of machine learning (ML) algorithms, in particular neural networks (NNs), have gained popularity for performing quantum parameter estimation [21, 27, 38, 43, 19, 9]. In contrast to Bayesian inference, NN based approaches provide faster inference time, owing to their highly parallelizable inference process, while retaining the benefits of likelihood-free methods like ABC as they can learn complex functions from input-output pairs of training examples due to their high expressive power [13]. Furthermore, the ABC-based estimation method presented in Ref. [10] requires the a priori preparation of multiple sets of measurement results for different values of $\theta$ to perform inference. While an ML approach also requires this for the initial model training phase it is typically no longer needed at inference time unlike the ABC method. In Ref. [43] it was shown that a NN can be trained to perform parameter estimation given input-output pairs of quantum measurement results from a continuous measurement of a quantum probe system and the corresponding parameter values which generated those measurement results. This approach however utilized a large amount of training data and only provides point estimates without uncertainty quantification. Here we extend this approach to provide uncertainty estimates, thereby retaining a benefit of Bayesian approaches, and also drastically reducing the amount of training data required to train the model. We also study the robustness of our method to different types of noise in the training data and measurement results provided to the model at inference time.
@@ III.1 Deep ensembles
Model ensembling is a simple but powerful method in machine learning. Ensemble methods employ multiple models, typically ‘weak learners’ as individually these models have suboptimal performance, to create a single overall model by aggregating the predictions of the component models. The resulting model often has better performance, in terms of both bias and variance, than the individual models in isolation [22]. A state-of-the-art method for predictive uncertainty quantification for deep learning based models are deep ensembles [31, 39, 17]. These models quantify epistemic (model) uncertainty, i.e. uncertainty that arises from the choice of model parameters, due to the use of of an ensemble of independent models. This type of uncertainty is reducible if one can procure more data. They also quantify aleatoric (data) uncertainty, i.e. uncertainty that is inherent to the process which produces the training data, by being trained to minimize a negative log likelihood loss function [17, 49, 47]. Unlike model uncertainty, data uncertainty is not reducible [49]. For regression problems, such as the one considered in this work, deep ensembles work by training $M$ independent NNs to learn a continuous probability distribution over predictions. The overall model is then treated as a mixture of these distributions. Model diversity emerges from the random initialization of each network at the start of training as well as the randomness from the mini-batch sampling that occurs during training via stochastic gradient descent.
Deep ensembles are preferable to Bayesian NNs (BNNs) [36], which are another common tool for quantifying uncertainty in deep learning, for two reasons. Firstly they require minimal changes to the standard NN training process and their training is easily scalable via parallelization since each network in the ensemble can be trained completely independently. Secondly BNNs face the same computational problem as Bayesian inference in general i.e. the intractability of computing the posterior distribution over model parameters exactly [36, 31]. In fact due to the large number of parameters in NNs standard Bayesian sampling methods like MCMC are often too computationally expensive to use for BNN training and methods such as variational inference, which are faster but have weaker theoretical guarantees compared to MCMC, must be used [17].
@@ III.2 Model training
respectively [31]. A useful property of Gaussian mixtures is that they are universal approximators for probability densities [20]. Note, however, that the log likelihood need not be that of a Gaussian, the negative log likelihood of any probability distribution may be used if deemed appropriate and one is able to compute the relevant moments.
An alternative approach to training an ensemble of $M$ NNs would be to simply train each network to minimize the mean squared error (MSE) loss function. The uncertainty estimates would then be computed as the empirical variance of the predictions made by the $M$ networks. While this approach is more straightforward it leads to the model producing highly overconfident uncertainty estimates as MSE only optimizes for predictive accuracy and not predictive uncertainty [31]. This highlights the crucial role played by selection of a proper loss function like the Gaussian negative log likelihood in training a deep ensemble to produce properly calibrated uncertainty estimates.
@@ III.3 Hyperparameter Tuning
Ref. [43] trained single NNs on a dataset of $4\times 10^{6}$ trajectory-parameter pairs [44]. With the use of hyperparameter tuning we are able to achieve the same level of performance with only $1\%$ ($4\times 10^{4}$) of the original training data. This is also much lower than the amount of data needed for ABC which used 1000 trajectories generated for 100 different values of $\Delta$ which results in a total dataset size of $10^{5}$ [10, 12]. Note also that the ABC approach used longer trajectories with $N=200$ time delay measurements as opposed to the $N=48$ used in this work. We split the training data into training and validation sets with a $80\%$ and $20\%$ split respectively. A commonly used approach to train ensembles models is the use of bootstrap aggregation (bagging) which consists of training ensemble members on different bootstrapped samples of the original dataset. However, this is mainly advantageous for ensemble models comprised of weak learners, such as tree-based ensemble models, and worsens performance when used to training ensembles of strong learners like deep neural networks [33, 31]. We therefore did not utilize bagging in our experiments.
We use the tree-structured Parzen estimator optimization algorithm provided by the Optuna framework to perform hyperparameter tuning of the single network models presented in this work [3]. We utilize both trail pruning and early stopping during our tuning and training processes respectively and tune for 20 trials. For the single models the loss function used during training was also tuned as part of the hyperparameter tuning process with the possible choices being the root mean squared error (RMSE) loss or the mean squared logarithmic error (MSLE) loss [43]. Further details on the tuned hyperparemeters can be found in Appendix B. For experiment tracking, validation metric logging, and model storage we utilize the open source framework MLflow [54]. We make the MLflow logs which contain learning curve plots, model weights, and hyperparameter settings for all the models trained in the present work available at [4].
III.4 Predictive accuracy preservation under ensembling
In this section we examine the performance of our ensemble approach based on predictive accuracy. This is crucial as we must ensure that the training of the ensemble does not sacrifice predictive accuracy in favor of well calibrated uncertainty estimates.
@@ III.5.2 Uncertainty quantification under data shift
(b) Average predicted variance $\sigma^{2}$ of deep ensemble for test data generated from the training distribution, which includes no time jitter noise, and test data with time jitter noise following the Gaussian distribution $\mathcal{N}(0,\sigma_{\tau})$.
@@
An important property to examine for models which provide uncertainty estimates is how the predicted uncertainty changes when the model is provided out-of-distribution examples e.g. in the case of data shift [31]. Ideally we would want our model to have higher predictive uncertainty when given input data from a distribution that is far from the training distribution i.e. we would expect the predicted Gaussian mixture to have greater variance $\sigma^{2}$. Fig. 7(a) shows the average predicted uncertainty of the deep ensemble model when given sets of time delays from systems in which the Rabi frequency differs from the value used to generate the training data. We can see that as the Rabi frequency deviates further from $\Omega=1$ the predictive uncertainty of our model increases.
This can be used to detect drift in the experimental data at inference time analogous to what has been demonstrated in Ref. [48] for image classification in industrial processes. For instance if one trains a deep ensemble on time delay data which were generated using a detector with minimal time jitter noise we would expect the ensemble, at inference time, to produce larger uncertainty estimates when presented with time delay measurements generated using a detector with time jitter noise. We can see that this is indeed the case in Fig. 7(b) which shows the average predicted uncertainty from an ensemble model, which was trained on time delay data with no jitter noise, for trajectories generated using varying levels of time jitter noise. This can be useful as a way of detecting imperfections or miscalibration in the experiential setup (in this case detector defectiveness) used to generate the time delay results given to the model at inference time.
IV Comparison to Approximate Bayesian Computation
Here we will compare the performance of our likelihood-free parameter inference method with the state-of-the-art likelihood-free method presented in Ref. [10] i.e. approximate Bayesian computation (ABC). For this comparison we consider the more complex optomechanical system studied in the same work.
IV.1 Quantum system
The system consists of an optical cavity which is driven by an external laser with one of the mirrors being free to oscillate which constitutes the mechanical mode of the system [11, 10]. The Hamiltonian of the closed system is given by
${}\hat{H}=-\Delta\hat{a}^{\dagger}\hat{a}+\omega_{M}\hat{b}^{\dagger}\hat{b}+\frac{\Omega}{2}(\hat{a}+\hat{a}^{\dagger})+g\hat{a}^{\dagger}\hat{a}(\hat{b}+\hat{b}^{\dagger})$ (9)
where $\hat{a}$ and $\hat{b}$ are the annihilation operators of the optical and mechanical modes respectively, $\Omega$ is the laser Rabi frequency, $g$ is the optomechanical coupling strength, and $\omega_{M}$ is the mechanical frequency. The mechanical mode interacts with a thermal bath by both emitting and absorbing phonons while the optical mode leaks photons which are continuously detected. The master equation for this open system is given by
${}\frac{\partial}{\partial t}\hat{\rho}(t)=$ $-i[\hat{H},\hat{\rho}(t)]+\kappa\left(\hat{a}\hat{\rho}(t)\hat{a}^{\dagger}-\frac{1}{2}\{\hat{a}^{\dagger}\hat{a},\hat{\rho}(t)\}\right)$ (10)
$+\gamma(\bar{m}+1)\left(\hat{b}\hat{\rho}(t)\hat{b}^{\dagger}-\frac{1}{2}\{\hat{b}^{\dagger}\hat{b},\hat{\rho}(t)\}\right)$ (11)
$+\gamma\bar{m}\left(\hat{b}^{\dagger}\hat{\rho}(t)\hat{b}-\frac{1}{2}\{\hat{b}\hat{b}^{\dagger},\hat{\rho}(t)\}\right)$ (12)
where $\bar{m}=(\exp{(\omega_{M}/k_{B}T)-1})^{-1}$ is the mean phonon number of the bath, $T$ is the temperature of the bath, $\kappa$ is the decay rate of the optical mode and $\gamma$ is the decay rate of the mechanical mode.
Specifically, we consider the nonlinear regime of the system which arises when the coupling strength between the optical and mechanical modes is strong [10] and therefore the analytic Gaussian formalism is no longer sufficient to describe the system dynamics [11]. In this case no analytic solution to the master equation is available [10], which precludes the derivation of a likelihood function, thus necessitating the use of likelihood-free methods such as ABC. Notably one can set the detuning value as $\Delta_{n}=-ng^{2}/\omega_{M}$, where $n\in\mathbb{Z}^{+}$, so that $n$-photon emissions become favored by the system resulting in photon bunching in the $n\geq 2$ regime [11, 10]. Note that the data used here has $g=4$ and $\omega_{M}=4\sqrt{2}$ [10].
IV.2 Predictive accuracy benchmark
We train our deep ensemble to again minimize the Gaussian negative log likelihood loss. For the training data we use the data library from Ref. [10]. The data contains time delay trajectories of $80$ photon detections for 100 evenly spaced values of the detuning $\Delta$ across the range [10,0][-10,0] with $4,000$ trajectories for each unique $\Delta$. These data were generated using parameters We randomly split the library using an 80/20 split between training and testing data and stratify by $\Delta$ value to ensure each set is balanced and that we use a uniform prior for the ABC algorithm as done in [10]. Due to the nature of the dynamics of the system in the nonlinear regime, i.e. the system can generally not be assumed to reset to a consistent state with each photon emission like in the case of the TLS, the assumption of independence and identical distribution of the time delays in a given trajectory no longer holds. While these additional time delay correlations allow for potential improvement in the ability to perform sensing tasks [11] they also lead to the smoothed histogram neural network layer used previously no longer being optimal.
Two tools that are more appropriate in this setting are recurrent neural network layers such as a long-short term memory layer or a one-dimensional convolutional layer. We use the latter as it is generally simpler to implement and more amenable to parallelization than recurrent architectures [51]. The specific architecture of the deep ensemble networks are shown in Table 1. We use an ensemble of $M=5$ networks with $\beta_{1}=0.87$, $\beta_{2}=0.93$, a learning rate of $10^{-3}$ for the Adam optimizer, and a batch size of $250$ over 50 epochs.
Table 1: Neural network architecture. The notation $M\rightarrow N$ denotes a mapping from $M$ features to NN features.
Module Configuration
Convolutional Layers
Conv1d $1\rightarrow 100$, kernel: 6, stride: 1, padding: 1
BatchNorm1d 100 features
ReLU activation
Conv1d $100\rightarrow 200$, kernel: 6, stride: 1, padding: 1
BatchNorm1d 200 features
ReLU activation
Dense Layers
Linear $200\rightarrow 128$
BatchNorm1d 128 features
ReLU activation
Linear $128\rightarrow 64$
BatchNorm1d 64 features
ReLU activation
Linear $64\rightarrow 32$
BatchNorm1d 32 features
ReLU activation
Output Layer
Linear $32\rightarrow 2$
When benchmarking the ABC method we treat the training data as the data library. While the pseudocode for the ABC algorithm is given in Ref. [10] we give a more detailed implementation in Algorithm 1 for completeness.
Algorithm 1 Approximate Bayesian computation algorithm
1:Library of trajectory-parameter pairs $\mathcal{D}$, number of times to sample from the library $\nu$, acceptance threshold for sample $\epsilon$, summary statistic function $S(\cdot)$, discrepancy measure function $\delta(\cdot)$, observed trajectory to perform parameter inference on $D$
2:Initialize empty histogram $H$ of $\theta$’s
[…]
4:while $i<\nu$ do
5:  Sample, without replacement, one trajectory-parameter pair $(D^{\prime},\theta)$ from $\mathcal{D}$
6:  if $\delta(S((D^{\prime},\theta)),S(D))\leq\epsilon$ then
7:   $H[\theta]\leftarrow H[\theta]+1$ $\triangleright$ Add one count to histogram bin corresponding to value $\theta$
8:  end if
9:  $i\leftarrow i+1$
10:end while
11:Normalize $H$
12:return $H$ $\triangleright$ Return the normalized histogram which serves as an approximation of the Bayesian posterior
We draw $\nu=10^{5}$ samples and $\epsilon=9$ as this was found to be more than enough for the method to converge for estimation of the same parameter and system in [10]. We also use the same histogram based summary statistic and $L_{2}$ discrepancy measure from [10] and similarly take the mean of the ABC posterior as our estimator of the parameter.
Figure 8: Performance of ABC and deep ensemble on the test set for the nonlinear optomechanical system.
Fig. 8 shows the performance of the two methods on our test set. We can see that for most values of $\Delta$ the deep ensemble achieves lower error than ABC and achieves competitive accuracy in all other cases. Interestingly, for the $\Delta_{n=3}$ regime we see that the ensemble greatly outperforms the ABC method suggesting it makes better use of the information encoded in the three-photon correlations present in the trajectory data.
V Computational Efficiency
@@
(b) Average inference times over 10 runs for HMC based and ABC based Bayesian inference and quantized deep ensemble for various numbers of trajectories.
@@
We also compare inference times by measuring the wall clock time of a trained quantized deep ensemble, the ABC method and an optimized likelihood-based Hamiltonian Monte Carlo (HMC) Bayesian inference routine for estimating $\Delta$ given delay measurements from different numbers of trajectories. We use the state-of-the-art No U-Turn Sampler (NUTS) provided by the PyMC framework as our specific HMC method [24, 2]. For inference with NUTS we use 4 chains and draw 1,000 posterior samples per chain which we parallelize over 4 cores. For inference with ABC we use the same implementation of the algorithm used in the previous section which leverages NumPy array vectorization when possible to ensure fair comparison. The average inference time for the three methods over 20 different runs of the inference on an AMD Ryzen 9 3900X 12-Core Processor are shown in Fig. 9(b). We can see that the highly parallelizable nature of neural network inference, i.e. batched inference, allows for orders of magnitude faster inference speed over the NUTS and ABC method despite the need to perform a forward pass for each of the 10 NNs in the deep ensemble. The ability to parallelize inference with a neural network, due to the fact that the network is carrying out simple linear algebraic operations, contrasts with the inherently sequential nature of MCMC based Bayesian inference due to the construction of the Markov chains.
VI Conclusion
We have proposed a method for using NNs to perform parameter estimation which provides uncertainty estimates in addition to point estimates of the parameters. We showed that these models provide good parameter estimates despite being optimized for both predictive accuracy and well calibrated uncertainty estimation. We also demonstrated our scheme is applicable to more complex systems in which the measurement results are not independent and identically distributed, unlike the case of the TLS studied in much of the present work, where recurrent or convolutional NNs become more appropriate to use over the permutation-invariant architectures discussed initially. Furthermore, our method can, in principle, be generalized to $n$-dimensional estimation problems by only adding an additional $O(n)$ neurons in the output layer of the network. We leave this for future work along with the exploration of other NN architectures, such as deep sets, that possess permutation invariance and the ability to naturally handle inputs of varying length [55]. In this work we performed hyperparameter tuning on a single model and used the optimal hyperparameters from that search to train the deep ensemble due to time constraints. An interesting question to answer would be if hyperparameter tuning on the ensemble level, while being more computationally expensive, could deliver even better results. Finally, in this work we considered ensembles in which all models have the same architecture but this is by no means a necessity. Further exploration of how using ensembles of networks with different architectures, e.g. different activation functions and numbers of hidden layers, would impact ensemble performance on the parameter estimation task would be useful, in particular with respect to how this affects the uncertainty estimates as this approach may better capture epistemic (model) uncertainty.